# JuliaQCD: Portable lattice QCD package in Julia language

### v1

@@ Abstract
We develop a new lattice gauge theory code set JuliaQCD using the Julia language. Julia is well-suited for integrating machine learning techniques and enables rapid prototyping and execution of algorithms for four dimensional QCD and other non-Abelian gauge theories. The code leverages LLVM for high-performance execution and supports MPI for parallel computations. Julia’s multiple dispatch provides a flexible and intuitive framework for development. The code implements existing algorithms such as Hybrid Monte Carlo (HMC), many color and flavor, supports lattice fermions, smearing techniques, and full QCD simulations. It is designed to run efficiently across various platforms, from laptops to supercomputers, allowing for seamless scalability. The code set is currently available on GitHub https://github.com/JuliaQCD.
@@ I Introduction
Quantum Chromo-dynamics (QCD) is the fundamental theory describing the interactions of quarks and gluons, the elementary particles that constitute the subatomic world. Since they can be created and annihilated from the vacuum through relativistic effects, they must be handled as a quantum many-body system with uncertain particle number. QCD is a quantum field theory and an accurate description of our universe.
However, perturbation theory fails to apply in QCD due to the strong coupling constant. Lattice QCD addresses this challenge by employing lattice regularization, introducing a discrete spacetime cutoff to regulate the path integral. This approach allows for the computation of quantum expectation values of physical observables.
Lattice QCD was first formulated by K. Wilson in 1974 Wilson (1974), and the numerical study of QCD using this method was initiated by M. Creutz Creutz (1980). This framework enables the detailed study of the QCD vacuum structure and provides precise insights into the Standard Model. Additionally, it facilitates investigations into phenomena such as the quark-gluon plasma and candidates for dark matter.
As we mentioned above, to formulate lattice QCD, a spacetime lattice cutoff is introduced. This cutoff is not merely an approximation for numerical calculations; it fundamentally defines QCD itself. Traditionally, quantum field theory is formulated on a continuous spacetime, but in the presence of interactions, such theories suffer from divergences. Although these divergences can be renormalized, they are inherent in the continuum formulation.
In QCD, we aim to calculate the quantum expectation value of a gauge-invariant operator $O$  [^1] For certain purposes, we sometimes calculate the expectation value of a gauge variant operator. In such cases, gauge fixing is required due to Elizur’s theorem Elitzur (1975).. In the path integral formalism, this is expressed as
@@
where $q$ and $\bar{q}$ are the quark and anti-quark fields, respectively, and $U$ is the gauge field on the lattice. The term $S_{\text{g}}[U]$ represents the gauge field action, and $S_{\text{f}}[U,q,\bar{q}]$ denotes the action for the quarks. The factor $Z$ is a normalization constant ensuring that $\langle 1\rangle=1$. One can obtain observables in QCD evaluating this integral.
@@
The standard algorithm used for simulating lattice QCD is the Hybrid Monte Carlo (HMC) algorithm Duane et al. (1987). HMC has become the de facto standard because of its efficiency and effectiveness in dealing with the large-dimensional integrals for lattice gauge fields with dynamical fermions. It has been successfully used in a wide range of studies, from calculating hadron spectrum Ali Khan et al. (2002) and form factors to determining the critical temperature ($T_{c}$) of QCD. The success of lattice QCD in these areas highlights the power of numerical methods in exploring non-perturbative aspects of quantum field theory.
@@
Recent developments in machine learning (ML) have opened up new avenues for enhancing lattice QCD calculations. Many people are developing new algorithms that integrate machine learning techniques into lattice QCD simulations. The importance of prototyping in this context cannot be overstated, as we aim to create efficient algorithms for lattice QCD that leverage ML techniques. One of the central trade-offs in this process is between code simplicity and execution speed. While ease of implementation is important, a code that is too slow is unusable in practice. Therefore, we focus on optimizing the total time required for lattice QCD studies, which includes both the time spent developing the algorithm and the time taken for its execution.
@@
Portability and parallelizability are also crucial considerations. The code must be able to run efficiently on a wide range of platforms, from laptops to supercomputers, and should be optimized to reduce the overall computational time. This allows researchers to take advantage of the available computational resources and scale their simulations as needed.
@@
The Julia language is a new open-source scientific programming language using LLVM Bezanson et al. (2012). Unlike interpreters, Julia is a compiled language with a just-in-time (JIT) compiler. This allows it to achieve performance comparable to Fortran and C jul , while maintaining the simplicity of Python. Julia’s flexibility and speed make it an ideal choice for developing new lattice QCD algorithms, as it allows for rapid prototyping without sacrificing performance. Thanks to these advantages, Julia is attractive language for high energy physics Eschle et al. (2023).
@@
In this paper we introduce a new lattice QCD code JuliaQCD written in Julia language. The motivation behind the development of this code set is to minimize the combined time of algorithm implementation and execution. Existing languages such as Python are too slow for large-scale lattice QCD simulations, while C++/C/Fortran present challenges in terms of ease of use and development time. Julia strikes the right balance, offering both the speed required for intensive computations and the simplicity needed for fast prototyping. Our goal is to create a flexible and high-performance framework that can be easily adapted to new developments in both lattice QCD and machine learning.
@@ I.1 JuliaQCD project
JuliaQCD is a suite of first open-source lattice QCD code by Julia programming language. (Ultimate) High performance is not within our scope but we are focus on prototyping with good scale. We developed this code with the following goals:
@@
The system offers excellent portability, as it runs smoothly on any machine with Julia installed, without the need for complicated setup processes. It is also quick and easy to set up, enabling users to start working in less than 5 minutes without requiring complex compilation steps. In addition, the system provides a comprehensive suite that supports both full-QCD and quenched configuration generation, along with a variety of measurement tools. Its highly modifiable structure makes it ideal for rapid prototyping and experimenting with different approaches. Finally, the system delivers competitive performance, comparable to the speed of traditional Fortran 90 codes, ensuring efficiency without sacrificing flexibility.
I.2 Related Works
In the broader landscape, traditional codes such as the MILC code and the Columbia Physics System (CPS) have also played significant roles. Despite their tradition, these codes remain functional and are used in various research contexts. MILC, in particular, has been a cornerstone for many researchers, supporting a wide range of computations, including those utilizing GPU architectures mil . CPS continues to be a relevant reference point for its C-based implementation. FermiQCD Di Pierro (2002) has contributed to the field. FermiQCD, a collection of C++ classes and parallel algorithms, facilitates fast development of lattice applications and includes optimizations for cluster computing, such as SSE2 instructions. Its focus on distributed processing has allowed it to remain an important tool in parallel lattice QCD.
@@ I.2 Related Works
The development of lattice QCD codes has been significantly influenced by contributions from various international communities, including notable efforts from Japan. Two key projects, the Lattice Tool Kit in Fortran 90 (LTK) Choe et al. (2002) and Bridge++ Akahoshi et al. (2022), originate from this community. LTK, primarily written in Fortran, has served as a foundational tool, though it is not optimized for contemporary high-performance computing (HPC) environments. In contrast, Bridge++ was developed focusing on providing a comprehensive and efficient toolset for lattice QCD simulations on modern supercomputers. Both projects reflect Japan’s strong engagement in advancing computational techniques for lattice QCD.
In recent years, there has been a trend towards utilizing C++, like Bridge++, for its advanced features and improved performance in complex simulations. Grid Boyle et al. (2015) is a prime example of this trend, offering robust support for both CPU and GPU computations. The Grid Python Toolkit (GPT) extends this functionality to Python, allowing for easy integration with data analysis and machine learning workflows. Another significant development is SIMULATeQCD Mazur et al. (2024), which specializes in multi-GPU implementations, enabling high-performance lattice QCD computations. Additionally, QUDA (QCD on CUDA) Clark et al. (2010) serves as a sub-library that provides specialized routines for running lattice QCD computations on NVIDIA GPUs. By leveraging CUDA architectures, QUDA offers substantial speedups for specific operations within larger QCD packages, rather than functioning as a standalone QCD software suite.
The increasing adoption of Python in scientific computing has led to the development of tools like Lyncs-API and pyQCD Spraggs (2014) and also GPT. These projects provide high-level interfaces for lattice QCD calculations, combining Python’s user-friendly syntax with the computational efficiency of C++ backends. Lyncs-API, as discussed by Bacchio et al. (2022), offers a flexible API for a variety of QCD computations, while pyQCD similarly integrates Python and C++ for efficient lattice QCD research.
This landscape illustrates the evolution of lattice QCD software from traditional Fortran and C implementations to modern, multi-language frameworks designed for the latest HPC environments. LatticeQCD.jl contributes to this ongoing evolution by leveraging Julia’s capabilities, providing a new, efficient, and accessible option for researchers.
@@ II Software description
We support lattice gauge theory in 4-dimensional Euclidean spacetime. We cover most of standard technologies. Also we implement a Wizard for parameter files. Our code LatticeQCD.jl has following functionalities.
@@
Fermion action
  • Wilson (2 flavor)
  • Staggered fermion (1-8 tastes)
  • Standard Domain-wall (2 flavor)
@@
  • Momentum projected pion correlator (Wilson, staggered)
@@
  • Topological charge (plaquette, clover, and O($a^{2}$) improved definition)
@@
  • Cold/Hot start for $\mathrm{SU}(N_{c})$. One instanton configuration for SU(2)
  • Heatbath for $\mathrm{SU}(N_{c})$ overrelaxation for a general gauge action
@@
  • HMC (2 flavor Wilson) with $\mathrm{SU}(N_{c})$ with a general gauge action
  • HMC (4 taste staggered fermions) with $\mathrm{SU}(N_{c})$ with a general gauge action
  • RHMC (any flavor staggered) with $\mathrm{SU}(N_{c})$ for a general gauge action
  • $\mathrm{SU}(N_{c})$ stout smeared dynamical fermions
@@
  • JLD format (Default binary file for Julia, one of HDF5)
  • Text file for Bridge++ (Bridgetext)
If one specified other than $N_{f}=4,8$ with the staggered fermion HMC, RHMC is automatically used. For a machine with the Apple Silicon, $N_{f}=1-8$ is available. Parallelization is supported by MPI, which can be used in Gaugefields.jl and LatticeDiracOperators.jl (see Appendix B.4). We perform consistency checks with several references. See Appendix E.
@@ II.1 Code Structure
JuliaQCD is a project name and it is constructed by several packages.
  • LatticeQCD.jl : LatticeQCD.jl is a wrapper of the following packages. Interface of this package is user friendly.
  • Wilsonloop.jl: Wilsonloop.jl helps us to treat the Wilson loops and generic Wilson lines in any $N_{c}$ and dimensions. Wilson lines can is defined symbolically.
  • Gaugefields.jl: Gaugefields.jl is a package for lattice $\mathrm{SU}(N_{c})$ gauge fields. It handles gauge fields (links), gauge actions with MPI, and autograd. This can generate quenched configurations.
  • LatticeDiracOperators.jl: LatticeDiracOperators.jl is a package for Dirac operators and fermions on the lattice. It handles pseudo-fermion fields with various lattice Dirac operators, fermion actions with MPI. This can generate configurations with dynamical fermions.
@@
Figure 1: Structure of whole packages (JuliaQCD).
@@
We show the example code to perform lattice QCD calculations using LatticeQCD.jl. One can start a lattice QCD calculation in 5 steps.
@@
    Install Julia language from Julialang.org.
@@
    In Julia REPL, press the ] key to enter the package mode and type,
@@
    and then press the “return” key. Press the “backspace” key (or “delete” key for Mac) to exit the package mode. One can get the latest version via add LatticeQCD#master. All dependencies will be solved automatically.
@@
    One will get results. One can write a parameter file.
@@ III.1 User interfaces
We support the following two user interfaces for LatticeQCD.jl,
@@
    General interface (Experience with another code, for batch jobs, customized purpose)
Usage 1 is explained above. For Usage 2, in Julia REPL, press the ] key to enter the package mode and type,
@@
Then, LatticeQCD.jl will be installed.
The “PARAMETER_FILE” can be created through the wizard. To use the wizard on the shell, please write the following code and save it as wizard.jl,
@@
Please write the following code and save it as run.jl,
@@
Then, one can execute it like this,
@@ IV Conclusion and Outlook
In this work, we introduce JuliaQCD, a new and flexible lattice QCD code set developed in the Julia programming language. By leveraging Julia’s high-performance just-in-time compilation and ease of use, this project aims to strike a balance between rapid prototyping and computational efficiency. JuliaQCD offers a range of functionalities, including support for full-QCD simulations and various smearing techniques, while maintaining scalability across different computational environments from personal systems to high-performance computing clusters.
@@
While the performance of JuliaQCD is comparable to established Fortran, its focus remains on flexibility and usability, providing a framework that can be easily extended to incorporate modern computational techniques. Future work will explore further optimizations, including deeper integration with GPU architectures and advanced machine learning techniques to further enhance the capabilities of lattice QCD simulations.

### v2

@@ Abstract
We present JuliaQCD, a new lattice gauge theory codebase developed in the Julia language. Julia is well suited to integrating machine-learning techniques and enables rapid prototyping and execution of algorithms for four-dimensional QCD and other non-Abelian gauge theories. The code leverages LLVM for high-performance execution and supports MPI for parallel computing. Julia’s multiple dispatch provides a flexible and intuitive framework for development. The code implements existing algorithms such as Hybrid Monte Carlo (HMC) and supports many colors and flavors, lattice fermions, smearing techniques, and full QCD simulations. It is designed to run efficiently across various platforms, from laptops to supercomputers, allowing seamless scalability. The codebase is currently available on GitHub at https://github.com/JuliaQCD.
@@ I Introduction
Quantum Chromodynamics (QCD) is the fundamental theory describing the interactions of quarks and gluons, the elementary particles that make up the subatomic world. Because they can be created and annihilated from the vacuum by relativistic processes, they must be treated as a quantum many-body system with an indeterminate particle number. QCD is a quantum field theory that accurately describes our universe.
However, perturbation theory fails in QCD due to the strong coupling. Lattice QCD addresses this challenge by employing lattice regularization, introducing a discrete spacetime cutoff to regulate the path integral. This approach enables the computation of quantum expectation values of physical observables.
Lattice QCD was first formulated by K. G. Wilson in 1974 Wilson (1974), and numerical studies of QCD using this method were initiated by M. Creutz Creutz (1980). This framework enables detailed studies of the QCD vacuum structure and provides precise insights into the Standard Model. Additionally, it facilitates investigations into phenomena such as the quark-gluon plasma and dark-matter candidates.
To formulate lattice QCD, a spacetime lattice cutoff is introduced. This cutoff is not merely a numerical approximation; it fundamentally defines QCD itself. Traditionally, quantum field theory is formulated on a continuous spacetime, but in the presence of interactions such theories suffer from divergences. Although these divergences can be renormalized, they are inherent in the continuum formulation.
In QCD, we aim to calculate the quantum expectation value of a gauge-invariant operator $O$  [^1] For certain purposes, we sometimes calculate the expectation value of a gauge-variant operator. In such cases, gauge fixing is required by Elitzur’s theorem Elitzur (1975).. In the path integral formalism, this is expressed as
@@
where $q$ and $\bar{q}$ are the quark and antiquark fields, respectively, and $U$ is the gauge field on the lattice. The term $S_{\text{g}}[U]$ represents the gauge field action, and $S_{\text{f}}[U,q,\bar{q}]$ denotes the action for the quarks. The factor $Z$ is a normalization constant ensuring that $\langle 1\rangle=1$. One can obtain observables in QCD by evaluating this integral.
@@
The standard algorithm used for simulating lattice QCD is the Hybrid Monte Carlo (HMC) algorithm Duane et al. (1987). HMC has become the de facto standard because of its efficiency and effectiveness in dealing with the high-dimensional integrals for lattice gauge fields with dynamical fermions. It has been successfully used in a wide range of studies, from calculating the hadron spectrum Ali Khan et al. (2002) and form factors to determining the critical temperature ($T_{c}$) of QCD. The success of lattice QCD in these areas highlights the power of numerical methods in exploring nonperturbative aspects of quantum field theory.
@@
Recent developments in machine learning (ML) have opened new avenues for enhancing lattice QCD calculations. Many researchers are developing new algorithms that integrate ML techniques into lattice QCD simulations. Prototyping is crucial in this context, as we aim to create efficient algorithms for lattice QCD that leverage ML techniques. One of the central trade-offs in this process is between code simplicity and execution speed. While ease of implementation is important, code that is too slow is unusable in practice. Therefore, we focus on optimizing the total time required for lattice QCD studies, which includes both the time spent developing the algorithm and the time required for its execution.
@@
Portability and parallelizability are also crucial considerations. The code must run efficiently on a wide range of platforms, from laptops to supercomputers, and should be optimized to reduce the overall computational time. This allows researchers to take advantage of the available computational resources and scale their simulations as needed.
@@
The Julia language is a new open-source scientific programming language built on LLVM Bezanson et al. (2012). Unlike interpreters, Julia is a compiled language with a just-in-time (JIT) compiler. This allows it to achieve performance comparable to Fortran and C jul , while maintaining the simplicity of Python. Julia’s flexibility and speed make it an ideal choice for developing new lattice QCD algorithms, as it allows for rapid prototyping without sacrificing performance. Thanks to these advantages, Julia is an attractive language for high-energy physics Eschle et al. (2023).
@@
In this paper we introduce a new lattice QCD code, JuliaQCD, written in the Julia language. The motivation behind the development of this codebase is to minimize the combined time of algorithm implementation and execution. Existing languages such as Python are too slow for large-scale lattice QCD simulations, while C++/C/Fortran present challenges in terms of ease of use and development time. Julia strikes the right balance, offering both the speed required for intensive computations and the simplicity needed for fast prototyping. Our goal is to create a flexible and high-performance framework that can be easily adapted to new developments in both lattice QCD and machine learning.
@@ I.1 JuliaQCD project
JuliaQCD is the first open-source lattice QCD code suite in the Julia programming language. (Ultimate) high performance is not within our scope, but we focus on prototyping with good scalability. We developed this code with the following goals:
@@
The system offers excellent portability, as it runs smoothly on any machine with Julia installed, without the need for complicated setup processes. It is also quick and easy to set up, enabling users to start working in less than 5 minutes without requiring complex compilation steps. In addition, the system provides a comprehensive suite that supports both full QCD and quenched configuration generation, along with a variety of measurement tools. Its highly modifiable structure makes it ideal for rapid prototyping and experimenting with different approaches. Finally, the system delivers competitive performance, comparable to traditional Fortran 90 codes, ensuring efficiency without sacrificing flexibility.
I.2 Related Work
In the broader landscape, traditional codes such as the MILC code and the Columbia Physics System (CPS) have also played significant roles. Despite their long history, these codes remain functional and are used in various research contexts. MILC, in particular, has been a cornerstone for many researchers, supporting a wide range of computations, including those utilizing GPU architectures mil . CPS continues to be a relevant reference point for its C-based implementation.
FermiQCD Di Pierro (2002) has also contributed to the field. FermiQCD, a collection of C++ classes and parallel algorithms, facilitates fast development of lattice applications and includes optimizations for cluster computing, such as SSE2 instructions. Its focus on distributed processing has allowed it to remain an important tool in parallel lattice QCD.
@@ I.2 Related Work
The development of lattice QCD codes has been significantly influenced by contributions from various international communities, including notable efforts from Japan. Two key projects, the Lattice Tool Kit in Fortran 90 (LTK) Choe et al. (2002) and Bridge++ Akahoshi et al. (2022), originate from this community. LTK, primarily written in Fortran, has served as a foundational tool, though it is not optimized for contemporary high-performance computing (HPC) environments. In contrast, Bridge++ was developed with a focus on providing a comprehensive and efficient toolset for lattice QCD simulations on modern supercomputers. Both projects reflect Japan’s strong engagement in advancing computational techniques for lattice QCD.
In recent years, there has been a trend toward utilizing C++, as in Bridge++, for its advanced features and improved performance in complex simulations. Grid Boyle et al. (2015) is a prime example of this trend, offering robust support for both CPU and GPU computations. The Grid Python Toolkit (GPT) extends this functionality to Python, allowing for easy integration with data analysis and machine learning workflows. Another significant development is SIMULATeQCD Mazur et al. (2024), which specializes in multi-GPU implementations, enabling high-performance lattice QCD computations. Additionally, QUDA (QCD on CUDA) Clark et al. (2010) serves as a sub-library that provides specialized routines for running lattice QCD computations on NVIDIA GPUs. By leveraging CUDA architectures, QUDA offers substantial speedups for specific operations within larger QCD packages, rather than functioning as a standalone QCD software suite.
The increasing adoption of Python in scientific computing has led to the development of tools such as Lyncs-API, pyQCD Spraggs (2014), and the Grid Python Toolkit (GPT). These projects provide high-level interfaces for lattice QCD calculations, combining Python’s user-friendly syntax with the computational efficiency of C++ backends. Lyncs-API, as discussed in Bacchio et al. (2022), offers a flexible API for a variety of QCD computations, while pyQCD similarly integrates Python and C++ for efficient lattice QCD research. This landscape illustrates the evolution of lattice QCD software from traditional Fortran and C implementations to modern, multi-language frameworks designed for contemporary HPC environments. LatticeQCD.jl contributes to this ongoing evolution by leveraging Julia’s capabilities, providing a new, efficient, and accessible option for researchers.
@@ II Software description
We support lattice gauge theory in 4-dimensional Euclidean spacetime. We cover most standard techniques. We also implement a wizard for parameter files. Our code, LatticeQCD.jl, has the following functionalities.
@@
Fermion actions
  • Wilson (2 flavors)
  • Staggered fermions (1-8 tastes)
  • Standard domain-wall (2 flavors)
@@
  • Momentum-projected pion correlator (Wilson, staggered)
@@
  • Topological charge (plaquette, clover, and $\mathcal{O}(a^{2})$-improved definition)
@@
  • Cold/Hot start for $\mathrm{SU}(N_{c})$. One instanton configuration for $\mathrm{SU}(2)$
  • Heatbath for $\mathrm{SU}(N_{c})$ & overrelaxation for a general gauge action
@@
  • HMC (2-flavor Wilson) for $\mathrm{SU}(N_{c})$ with a general gauge action
  • HMC (4-taste staggered fermions) for $\mathrm{SU}(N_{c})$ with a general gauge action
  • RHMC (any-flavor staggered) for $\mathrm{SU}(N_{c})$ with a general gauge action
  • Stout-smeared dynamical fermions for $\mathrm{SU}(N_{c})$
@@
  • JLD format (Julia’s default binary file, HDF5-based)
  • Bridge++ text file (Bridgetext)
If one specifies other than $N_{f}=4,8$ with the staggered-fermion HMC, RHMC is automatically used. For a machine with Apple Silicon, $N_{f}=1-8$ is available. Parallelization is supported by MPI, which can be used with Gaugefields.jl and LatticeDiracOperators.jl (see Appendix B.4). We perform consistency checks against several references. See Appendix E.
@@ II.1 Code Structure
JuliaQCD is the project name, and it is composed of several packages.
  • LatticeQCD.jl : LatticeQCD.jl is a wrapper of the following packages. The interface of this package is user-friendly.
  • Wilsonloop.jl: Wilsonloop.jl helps handle Wilson loops and generic Wilson lines in any $N_{c}$ and dimensions. Wilson lines can be defined symbolically.
  • Gaugefields.jl: Gaugefields.jl is a package for lattice $\mathrm{SU}(N_{c})$ gauge fields. It handles gauge fields (links), gauge actions with MPI, and autograd. It can generate quenched configurations.
  • LatticeDiracOperators.jl: LatticeDiracOperators.jl is a package for Dirac operators and fermions on the lattice. It handles pseudofermion fields with various lattice Dirac operators, and fermion actions with MPI. It can generate configurations with dynamical fermions.
@@
Figure 1: Structure of the packages (JuliaQCD).
@@
We present example code to perform lattice QCD calculations using LatticeQCD.jl. One can start a lattice QCD calculation in 5 steps.
@@
    Install the Julia language from JuliaLang.org.
@@
    In Julia REPL, press the ] key to enter the package mode and type:
@@
    then press the “return” key. Press the “backspace” key (or “delete” key on Mac) to exit package mode. The latest version is available via add LatticeQCD#master. All dependencies will be resolved automatically.
@@
    One will obtain results. One can write a parameter file.
@@ III.1 User interfaces
We support the following two user interfaces for LatticeQCD.jl:
@@
    General interface (experience with another code, for batch jobs, customized purposes)
Usage 1 is explained above.
For Usage 2, in Julia REPL, press the ] key to enter the package mode and type:
@@
Then, LatticeQCD.jl will be installed. The “PARAMETER_FILE” can be created through the wizard. To use the wizard in the shell, please write the following code and save it as wizard.jl:
@@
Please write the following code and save it as run.jl:
@@
Then, one can execute it like this:
@@ IV Conclusion and Outlook
In this work, we introduce JuliaQCD, a new and flexible lattice QCD codebase developed in the Julia programming language. By leveraging Julia’s high-performance just-in-time compilation and ease of use, this project aims to strike a balance between rapid prototyping and computational efficiency. JuliaQCD offers a range of functionalities, including support for full QCD simulations and various smearing techniques, while maintaining scalability across different computational environments from personal systems to high-performance computing clusters.
@@
While the performance of JuliaQCD is comparable to established Fortran codes, its focus remains on flexibility and usability, providing a framework that can be easily extended to incorporate modern computational techniques. Future work will explore further optimizations, including deeper integration with GPU architectures and advanced machine learning techniques to further enhance the capabilities of lattice QCD simulations.